Using the method of substitution to help solve for this equation.
x+3y=1
-2x+2y=6
I am stuck this is what I have so far
x+3(-2x+6)=1
x-6+24=1
Am I on the right track?

Answer by jim_thompson5910(35256)

First solve -2x+2y=6 for y.

-2x+2y=6

2y=6+2x

2y = 2x+6

y = x+3 ... divide every term by 2

-------------------------------------------------------

Now turn to x+3y=1


x+3y=1

x+3(x+3)=1 ... plug in y = x+3

x+3x+9 = 1

4x+9 = 1

4x = 1-9

4x = -8


I'll let you finish solving for x. Once you have the value of x, use it in y = x+3 to find y.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------